Friday, 21 August

Friday, 21 August

Plenary Lectures

9:30 – 10:30  Auditorium Maximum H 105

Helmut H.W. Hofer, Courant Institute, New York University, New York, USA Dynamics, Topology, and Holomorphic Curves (I: p.8)

Chair: Vladimir Arnold, Steklov Mathematical Institute, Moscow, Russia

11:00 – 12:00  Auditorium Maximum H 105

Vladimir Voevodsky, Northwestern University, Evanston, USA Homotopy Theory of Algebraic Varieties (I: p.28)

Chair: Eric Mark Friedlander, Northwestern University, Evanston, USA

Invited Lectures

\underline{14:00 – 14:45}

04 Mikhail M. Kapranov, Northwestern University, Evanston, USA Operads and algebraic geometry (I: p.65) Chair: Enrico Arbarello

MA 005

05 Simon K. Donaldson, Stanford University, Stanford, USA Positive line bundles in symplectic geometry (I: p.69) Chair: Ulrich Pinkall

MA 001

06 Alexander N. Dranishnikov, University of Florida, Gainesville, USA and Steklov Math. Institute, Moscow, Russia Dimension theory and large Riemannian manifolds (I: p.79) Chair: Alexander S. Mishchenko

MA 004

09 Shuhei Hayashi, Waseda University, Tokyo, Japan Hyperbolicity, stability and the creation of homoclinic points (I: p.113) Chair: Jack Carr

H 1012

11 Fedor Smirnov, CNRS, LPTHE, Université Pierre et Marie Curie, Paris, France Space of local fields in integrable field theory and deformed Abelian integrals (I: p.137) Chair: Andreas Fring

H 1028

14 Emo Welzl, ETH Zürich, Zürich, Switzerland Halving Point Sets (I: p.175) Chair: Helmut Alt

H 2032

\underline{15:00 – 15:45}

04 Victor Vadimovich Batyrev, Universität Tübingen, Tübingen, Germany Mirror symmetry and toric geometry (I: p.60) Chair: Yujiro Kawamata

MA 005

05 François Labourie, Université Paris-Sud, Orsay, France Large Group Actions on Manifolds (I: p.73) Chair: Ulrich Pinkall

MA 001

06 Michael Hartley Freedman, Microsoft Corporation, Redmond, USA Topological views of computational complexity (I: p.80) Chair: Robert D. Edwards

MA 004

09 Michael Herman, CNRS, Université Denis Diderot / Paris 7, Paris, France Some open problems in dynamical systems (I: p.114) Chair: Thomas Bartsch H 1012 11 Roberto Schonmann, UCLA, Los Angeles, USA Metastability and the Ising model (I: p.136) Chair: Christian Borgs

H 1028

14 Miklós Ajtai, IBM Almaden Research Center, San Jose, USA Worst-case Complexity, Average-case Complexity, and Lattice Problems (I: p.165) Chair: Helmut Alt

H 2032

\underline{16:15 – 17:00}

02 William M. Kantor, University of Oregon, Eugene, USA Simple Groups in Computational Group Theory (I: p.41) Chair: Charles C. Sims

MA 005

07 Vera Serganova, University of California, Berkeley, USA On representations of Lie superalgebras (I: p.95) Chair: Olivier Mathieu

MA 004

08 Michael Christ, University of California, Berkeley, USA Singularity and Regularity — Local and Global (I: p.97) Chair: Daryl Geller

MA 001

10 Gustavo Alberto Ponce, University of California, Santa Barbara, USA On nonlinear dispersive equations (I: p.125) Chair: Mikhail Shubin

H 1012

12 Mark Freidlin, University of Maryland, College Park, USA Random and Deterministic Perturbations of Nonlinear Oscillators (I: p.141) Chair: Jürgen Gärtner

H 1028

16 & 17 William J. Cook, Rice University, Houston, USA The Traveling Salesman Problem (I: p.186) Chair: Günter M. Ziegler

H 2032

\underline{17:15 – 18:00}

02 Aleksandr Pukhlikov, Institute for Systems Analysis, Moscow, Russia Birational automorphisms of algebraic varieties (I: p.42) Chair: Victor Vadimovich Batyrev

MA 005

07 Shahar Mozes, Hebrew University, Jerusalem, Israel Products of trees, lattices and simple groups (I: p.93) Chair: Peter Sarnak

MA 004

08 Hart F. Smith, University of Washington, Seattle, USA Wave Equations with Low Regularity Coefficients (I: p.107) Chair: Gerald B. Folland

MA 001

10 Frédéric Hélein, CMLA, ENS-Cachan, Cachan, France Phenomena of compensation and estimates for partial differential equations (I: p.121) Chair: Gunther A. Uhlmann

H 1012

12 Jean-François Le Gall, DMI — École Normale Supérieure, Paris, France Branching processes, random trees and superprocesses (I: p.147) Chair: Donald Dawson

H 1028

17 Jan C. Willems, University of Groningen, Groningen, The Netherlands Open Dynamical Systems and their Control (I: p.195) Chair: Fritz Colonius H 2032

Short Communications and Poster Sessions

Short Communications: Friday, 21 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 01

Categories: Logic and Topology, Chair: Gerhard Preuß

MA 042 Michel Hebert Syntactic view on locally generated categories: abstract and concrete cases (p.7)

Dieter Leseberg

Filterconvergence as a convenient concept for Topology (p.8)

Sect.: 02

Groups II, Chair: Mikhail H. Klin

MA 311 Hubert Kiechle Frobenius groups with many involutions

(p.20)

Anatolii S. Kondratiev

Finite linear groups of small degree (p.20)

Wujie Shi

A new classification of groups (p.28)

Sect.: 07

Quantum Groups and Vertex Operator Algebras, Chair: Gerald Höhn

H 112 Kailash C. Misra Demazure Crystals for quantum affine Lie algebras (p.108)

Naihuan Jing

Quantized Kac-Moody algebras and symmetric functions (p.106)

Gerald Höhn

Vertex Operator Algebras and 3-Dimensional Topological Quantum Field Theories (p.105)

Sect.: 08

Univalent Functions I, Chair: Ruben A. Hidalgo

MA 041 Bruce O’Neill Boundary Behavior of Conformal Mappings Associated with Singular Measures

(p.145)

Teodor Bulboacă

Differential subordinations and univalent integral operators (p.119)

Om Parkash Ahuja

On Meromorphic Starlike and Convex Functions (p.113)

Sect.: 08

Integral Transforms II, Chair: Kun Soo Chang

MA 309 Yukio Kuribayashi On pseudofourier transforms (p.138)

Rama Rawat

Pompeiu problem and L^{p} analogues of Wiener-Tauberian for certain homogeneous spaces (p.151)

Kun Soo Chang

Integral transform and convolution of analytic funtionals on abstract Wiener space

(p.120)

Sect.: 08

Banach Spaces, Chair: David Yost

MA 043 Parviz Azimi A class of hereditarily complementably ℓ_{p} Banach spaces (p.115)

Robb C. Fry

Smooth Bump Functions on Banach Spaces (p.124)

David Yost

Twisted sums of classical Banach spaces (p.165)

Sect.: 10

Elliptic Equations II (Singularities & Index Theory), Chair: Vladimir S. Rabinovich

H 111 Vladimir Borisovich Vasil’ev Wave factorization of elliptic symbols

(p.223)

Igor I. Skrypnik

Degree of densely defined (S)_{+} operators and applications (p.222)

Vladimir S. Rabinovich

Elliptic boundary value problems in domains with non-Lipschitz singular points

(p.218)

Sect.: 10

Hyperbolic Equations I (Integrable Equations), Chair: Jiahong Wu

H 110 Igor Anders Long-time asymptotic behaviour of nondecaying solutions of KP-type equations

(p.198)

Tariel Kiguradze

On periodic solutions of nonlinear second order hyperbolic equations (p.209)

Jiahong Wu

Zero dissipation limits of wave equations (p.225)

Sect.: 12

Stochastic Processes I, Chair: Michael K.R. Scheutzow

H 107 Yuriy V. Kolomiets On behavior of solutions of evolution equations perturbed by random processes

(p.257)

Vassily A. Yaroshevsky

Estimation of large deviations of dynamic system phase (p.273)

Anna V. Naydonova

Analysis of Stochastic Recurrent Algorithms with the Random Reaction Time and Applications for Some Queueing Models (p.263)

Sect.: 12

Bayesian Methods, Chair: Michael Nussbaum

H 1029 Ammar Sarhan Empirical Bayes Estimators of Indexed Parameters of Exponential Families Olaf Bunke Asymptotic behauviour of Bayes estimates under possibly incorrect models

Sect.: 16

Mathematical Models in Biology, Chair: Satoshi Takahashi

H 106 Andrej Yu. Dubinsky Mathematical Modelling of electron and proton Transport in Chloroplasts (p.322)

Miguel Lara-Aparicio

Some mathematical aspects of the simulation of the motor circadian rhythm in crayfish (p.327)

Satoshi Takahashi

The oscillation of lateral asymmetry in scale-eating fish (p.333)

Sect.: 17

Optimality Conditions I, Chair: Bernd Kugelmann

H 2033 Vassili Dikoussar Optimal Control with Control-State Constraints (p.339)

Olga O. Vasilieva

Optimality Conditions of Singular Controls (p.354)

Sect.: 17

Variational Inequalities/Fractal Graphs, Chair: Walter Gomez Bofill

H 2013 Walter Gomez Bofill On the regularity of one-parametric variational inequalities (p.341)

Olga A. Zinchenko

The extreme graphs problems (p.356)

Anatolii Ivanovich Perov

Discrete matrix Lur’e equation (singular case) (p.347)

Short Communications: Friday, 21 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 01

Set Theoretic Topology, Chair: Dieter Leseberg

MA 042 Gerhard Preuß Convenient Topology — A New Branch of Topology (p.9)

Harald Nusser

Completions of probabilistic semiuniform convergence spaces (p.8)

Sect.: 02

Groups III, Chair: Andrew J. Woldar

MA 311 Vahagn H. Mikaelian Embeddings of countable groups into 2generator groups with additional properties (p.31)

Igor Ya. Subbotin

On groups with many periodic factorgroups (p.30)

Leonid A. Kurdachenko

On some infinite dimensional linear groups (p.21)

Sect.: 07

Quantum Groups and Poisson Homogeneous Spaces, Chair: Kailash C. Misra

H 112 Pu Zhang Ringel-Hall algebras and quantum groups

(p.111)

Ho Hai Phung

Rational representations of matrix quantum groups of type A_{n} (p.109)

Eugene A. Karolinsky

A classification of Poisson homogeneous spaces of complex reductive Poisson-Lie groups (p.106)

Sect.: 08

Quasiconformal Maps, Chair: Bruce O’Neill

MA 041 Alexander Jur’evich Vasil’ev Two point distortion theorems for quasiconformal mappings (p.162)

Evsei M. Dynkin

Smoothness of quasiconformal mappings at a point (p.122)

Vo Dang Thao

Area Problem for quasiconformal mappings (p.160)

Sect.: 08

Sobolev Classes and Matrix Functions, Chair: Ilya M. Spitkovsky

MA 309 Serguei Vodop’yanov On open and discrete mappings of Sobolev classes with integrable dilatation (p.162)

Vladimir V. Peller

Regularity conditions for vectorial stationary processes (p.147)

Ilya M. Spitkovsky

Almost periodic factorization and its applications (p.157)

Sect.: 08

C*-Algebras, Chair: George A. Elliott

MA 043 Martin Mathieu A simple local multiplier algebra (p.140)

Michael Frank

Conditional expectations of finite index and discrete group actions (p.124)

George A. Elliott

An abstract Weyl–von Neumann theorem (p.123)

Sect.: 10

PDE and Mathematical Physics, Chair: Roland Potthast

H 111 Johannes Elschner Existence, uniqueness, and regularity for solutions of the conical diffraction problem

(p.203)

Frank Jochmann

Existence and asymptotic behavior of weak solutions of the drift diffusion equations for semiconductors (p.207)

Roland Potthast

On mixed reciprocity relations in inverse scattering (p.217)

Sect.: 10

Navier-Stokes Equations, Chair: Takeyuki Nagasawa

H 110 Mikhail E. Bogovskii The problem of global smooth solutions for multidimensional Navier-Stokes equations solved (p.199)

Soichiro Katayama

Some semilinear wave equations with nonlinear damping terms (p.208)

Takeyuki Nagasawa

A new energy inequality and partial regularity for weak solutions of Navier-Stokes equations (p.215)

Sect.: 12

Stochastic Processes II, Chair: Peter Imkeller

H 107 Andrzej Makagon Periodically Correlated Processes and Mackey’s Imprimitivity Theorem (p.259)

Boris G. Martchenko

Concerning on a theorem for periodic in Slutskiy sense linear random processes

(p.260)

Andrew Ya. Olenko

On some properties of correlation and spectral functions (p.264)

Sect.: 12

Stochastic Algorithms, Chair: Peter Mathé

H 1029 Grigori Milstein The probability approach to numerical solution of nonlinear parabolic equations

(p.261)

Vadim I. Arkin

Optimal stopping problems with restrictions (p.248)

Peter Mathé

Efficiency of Monte Carlo Integration Using Markov Chains (p.260)

Sect.: 16

Game Theory and Mathematical Economics, Chair: Paul J. Campbell

H 106 Fausto Mignanego Rationality of players and credibility of strategies in leader-follower games (p.330)

Serguei Z. Levendorski

Generalizations of the Black-Scholes Formula for a Class of Truncated Lévy Processes (p.327)

Paul J. Campbell

Klondike solitaire (p.321)

Sect.: 17

Controllability and Nonlinear Systems, Chair: Shanjian Tang

H 2033 Serikbai A. Aisagaliev Theory and constructive methods of solving the boundary problems of controllability processes (p.335)

Zbigniew J. Bartosiewicz

Local observability, ideals, and finiteness conditions (p.336)

Shanjian Tang

Brockett’s problem of classification of finite-dimensional estimation algebras for Nonlinear Filtering Systems (p.352)

Sect.: 17

Multiobjective Optimization, Chair: Jan-J. Ruckmann

H 2013 Jürgen Guddat New Embeddings for Nonlinear Multiobjective Optimization Problems I (p.342)

Stefan M. Mititelu

Optimality and duality in nonsmooth vector programming (p.345)

Short Communications: Friday, 21 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 03

Algebraic Geometry I, Chair: Jürgen Wolfart

MA 311 Emmanuel Peyre Heights and Tamagawa measures (p.44)

Valentin E. Voskresenskii

Stably rational algebraic tori (p.46)

Jürgen Wolfart

Riemann surfaces with many automorphisms (p.47)

Sect.: 04

D-Modules/Vector Bundles I, Chair: Sabir M. Gusein-Zade

MA 042 Aleksandr G. Aleksandrov Gauß-Manin-Zusammenhang und logarithmische Picard-Fuchs-Systeme (p.49)

Vladimir Bavula

The Simple Holonomic Modules over the Ring of Differential Operators on a Smooth Algebraic Surface (p.50)

Ushadevi N. Bhosle

Moduli spaces of torsion-free sheaves on nodal curves (p.51)

Sect.: 06

Geometric Topology, Chair: Alexander N. Dranishnikov

H 112 Dusan Repovs On Banach-Mazur compacta and their geometric properties (p.96)

Washington Mio

On the geometric topology of homology manifolds (p.94)

Arkady Skopenkov

A Generalization of The Haefliger-Weber Theorem on Embeddings of Manifolds in ℝ^{m} (p.97)

Sect.: 09

Stability Theory of FDE, Chair: Géza Makay

H 111 Meirong Zhang The Fučik Spectrum and its Generalization for Asymptotically Positively Homogeneous Systems of Differential Equations

(p.197)

Mark R. Liberzon

Inners and Absolute Stability of Time-Lag Systems (p.181)

Sect.: 09

Dynamical Properties of Mechanical Systems, Chair: Dmitriy A. Chebanov

H 110 Daniel Massart On the convexity of the action of a Lagrangian (p.182)

Natalia Valeryevna Tkachenko

About some new solutions in a problem about motion of a body in a magnetic field

(p.193)

Dmitriy A. Chebanov

New exact solutions of the problem about motion of Lagrange’s gyroscopes system

(p.173)

Sect.: 11

Spectral Analysis I, Chair: George D. Raikov

H 107 Hermann Schulz-Baldes Quantum dynamics in presence of fractal spectra (p.245)

Michael Demuth

Stable absolutely and singularly continuous spectra due to integral conditions (p.232)

George D. Raikov

Eigenvalue asymptotics for the Dirac operator in strong constant magnetic fields

(p.243)

Sect.: 11

Thermodynamcis, Hydrodynamcis, Non-linear Dynamics I, Chair: Volodymyr Kotlyarov

H 1029 Xinyu He A probabilistic method for vortex motion in turbulent fluids (p.235)

Vladyslav V. Danishevs’kyy

A New Asymptotic Technique for Nonlinear Dynamic Boundary Value Problems

(p.231)

Volodymyr Kotlyarov

Scattering problem for compatible ordinary differential equations and nonlinear Schrödinger equation on the semi-axis

(p.238)

Sect.: 13

Paths and Cycles, Chair: Stephan Brandt

H 106 Angeles Carmona On the Shortest Paths in a Distance Regular Graph (p.275)

Min Aung

A lower bound for the circumference of a graph in terms of connectivity and minimum degree (p.283)

Henda Swart

Generalised Measures of Distance in Graphs (p.286)

Sect.: 14

Mathematical Aspects of Computer Science III, Chair: Clemens Gröpl

MA 043 Barbu Gheorghe A method for Computer Generation of Random Vectors (p.289)

Sven Hartmann

New results on the consistency of cardinality constraints (p.292)

Stefan Felsner

The Linear-Extension-Diameter (p.291)

Sect.: 15

Wavelets III (organized by Andreas Rathsfeld), Chair: Pierre Gilles Lemarié-Rieusset

H 2033 Siegfried Prößdorf Boundary element methods using multiwavelets (p.310)

Jörg Schult

Stability condition for spline and multiwavelet collocation (p.312)

Marie Farge

How to compute and model twodimensional turbulent flows using wavelets?

Short Communications: Friday, 21 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 03

Modular Forms, Chair: Yoichi Uetake

MA 311 Werner Hoffmann Spectral estimates for noncompact quotients (p.38)

Wladimir de Azevedo Pribitkin

Bol’s identity, fractional differentiation, and automorphic integrals (p.45)

Yoichi Uetake

A weak resolvent representation of a scattering matrix for automorphic functions

(p.46)

Sect.: 04

Deformations and Degenerations, Chair: Aleksandr G. Aleksandrov

MA 042 Kimio Miyajima CR construction of the semi-universal family of normal isolated singularities (p.57)

Eduardo S. Esteves

Degenerations of Weierstrass points (p.52)

Diego Paccagnan

Maximum number of cusps on algebraic plane curves and nodes on surfaces (p.59)

Sect.: 06

Hyperbolic 3-Manifolds, Chair: Nikolai Gusevskii

H 112 Vitalii S. Makarov Regular maps and hyperbolic 3-manifolds

(p.92)

Alexander D. Mednykh

The volumes and isometries of Heegaard genus two hyperbolic three manifolds

(p.93)

Andrei Yu. Vesnin

On 3-manifolds with cyclic symmetry (p.100)

Sect.: 09

Invariant Measures and Random Dynamics II, Chair: Gerhard Keller

H 111 Raphael Albrecht Periodic data, intersection types, and conjugacy of profinite surface automorphisms

(p.169)

Roman Hric

Chaos and Topological Entropy in Onedimensional Dynamics (p.177)

Gerhard Keller

Spectral stability of transfer operators under non-small perturbations (p.178)

Sect.: 09

Multifrequence Systems and Invariant Tori, Chair: Anatolij M. Samoilenko

H 110 Chong-Qing Cheng Aboundance of Lower Dimensional Invariant Tori in Hamiltonian Systems (p.174)

Miklós Rontó

A numerical-analytic method for parametrized BVPs (p.188)

Anatolij M. Samoilenko

Investigation of m-frequency oscillatory systems (p.189)

Sect.: 11

Spectral Analysis II, Chair: Andrey V. Tchourkin

H 107 Mark M. Malamud Inverse spectral analysis for first order systems (p.240)

Richard L. Hall

Geometric spectral inversion (p.235)

Andrey V. Tchourkin

Quantum stochastic model of gravitational wave interferometer detector (p.246)

Sect.: 14

Mathematical Aspects of Computer Science IV, Chair: Stefan Felsner

MA 043 Svetlana Asmuss On optimal algorithms of approximation under imprecise information (p.289)

Clemens Gröpl

Random OBDDs (p.292)

Andrew Adamatzky

On computation universality of excitable lattices

Sect.: 15

Partial Differential Equations III, Chair: Jianping Zhu

H 2033 P. Sobolevskii The “counter-example” to W. Littman’s counter-example of L_{p}-energetical inequality for wave equation (p.314)

Ilja Schmelzer

Grid Generation with Contravariant Geometry Description (p.311)

Sect.: 16

Interaction of Non-linear Waves, Chair: Brian T. Hayes

H 106 Motlatsi Pius Ramollo On the motion of a bore across a barrier

(p.331)

Saadet Erbay

Nonlinear interaction between long and short waves in a generalized elastic solid

(p.323)

Brian T. Hayes

Stress-Controlled and Velocity-Controlled Shear Waves in a Saturated Granular Medium (p.323)

Poster Session: Friday, 21 August, 14:00–15:45

Sect.: 08  Analysis

MA 144 Jelena Ausekle Compactness of Operators Acting in the Lorentz and Orlicz Sequence Spaces

(p.115)

Prodipeswar Bhattacharyya

On the domain of normality in iteration theory (p.117)

Davor Butkovic

Transformations of regular almost convergent sequences (p.119)

Rainis Haller M(r, s)-inequality and reflexivity (p.128)

Katsunori Iwasaki

Polyhedral harmonic functions and invariant theory (p.132)

Volodymyr M. Kadets

The Lyapunov property of certain Banach spaces (p.133)

Shuichi Ohno Topological structure of the space of composition operators on H^{∞} (p.145)

Eve Oja

Lifting approximation properties and ideals of compact and approximable operators

(p.145)

Zdzislaw Otachel

Cone Preorders in Measurable Function Spaces (p.145)

Märt Põldvere On subspaces of Banach spaces having Phelps’ uniqueness property (p.148)

Sen-Yen Shaw

Convergence rates of regularized approximation processes (p.156)

Panagiotis D. Stratigos

On general measure repleteness and completeness (p.158)

Sect.: 13

Combinatorics

MA 144 André Kündgen Turán problems for weighted graphs

(p.280)

Frank H. Lutz

Combinatorial manifolds with vertextransitive automorphism group (p.281)

Slobodan K. Simić

On graphs whose second largest eigenvalue does not exceed the golden section (p.284)

Indulis Strazdins Universal classification of Boolean functions by the affine and linear groups

(p.285)

Kai-Yao Xiang

The Minimal Group with Graph and Its Permutation Elements (p.287)

Tomoyuki Yoshida

Exponential formulas and Krull-Schmidt categories (p.288)

Poster Session: Friday, 21 August, 16:15–18:00

Sect.: 05

Differential Geometry and Global Analysis

MA 144 Philip P. Boalch Symplectic geometry and isomonodromic deformations (p.65)

Ilia A. Bogaevski

Singularities of convex hulls of closed hypersurfaces (p.65)

Cristina Draper

Algebras related to Lie triple systems and applications to differential geometry

(p.67)

Birte Feix Hyperkähler metrics on cotangent bundles

(p.68)

Lew Friedland

Spectral geometry on almost Hermitian manifolds (p.69)

Vladimir A. Geyler

Periodic point interactions on a Riemannian manifold (p.251)

Valery Konnov Differential geometry of algebraic manifolds (p.73)

Alexei G. Kovalev

Gluing theorems for anti-self-dual metrics (p.73)

Vasily Olegovich Manturov

Vertical Atoms, Chord Diagrams, and Knots (p.76)

Stana Nikcevic Curvature Homogeneous Spaces and Related Topics (p.77)

Marc Pauly

Singular monopoles on compact 3manifolds (p.77)

Vyacheslav D. Sedykh

On classes of convex curves in projective spaces (p.80)

Mamuka Shubladze The topology of hypersurface singularities with codimension one (p.80)

Peter Topalov

A family of 3-integrable metrics (p.83)

Paweł Walczak

Foliations, finitely generated groups, and pseudogroups of transformations of manifolds carry some dynamics (p.83)

Zuoling Zhou The Hausdorff measures of Sierpinski gasket and Koch curve (p.84)

Special Activities

\underline{14:00 – 18:00 H 104}

Mathematical Software Chair: Winfried Neun, ZIB, Berlin, Germany In this session, a broad spectrum of mathematical software including symbolic and numerical systems as well as interfaces and problem solutions will be presented. Details of the presentations can be found on the ICM’98 web server at URL http://elib.zib.de/ICM98/MathSoftware .

14:00 – 14:30 John H. Maddocks and Randy C. Paffenroth, EPF Lausanne, Switzerland VBM

14:40 – 15:10 Horst G. Zimmer et. al., Universität Saarbrücken, Germany SIMATH

15:15 – 15:45 Karim Belabas, Université Paris–Sud, Orsay, France PARI / GP

15:45 – 16:15 Break

16:15 – 16:45 Christian Wieners et. al., Universität Stuttgart, Germany UG

16:50 – 17:20 Elias Deeba, University of Houston–Downtown, Texas, USA Learning Linear Algebra Using an Interactive Text

17:30 – 18:00 David M. Gay and Robert Fourer, Murray Hill, N.J., USA AMPL \underline{19:30 – 21:00 H 104}

Events and policies: Effects on women in mathematics Panel discussion organized by women from the Association for Women in Mathematics (AWM), the European Women in Mathematics (EWM), and the Committee on Women and Mathematics of the European Mathematical Society

\underline{21:15 – 22:00 H 104}

Women and mathematics across cultures Film, directed by Marjatta Näätänen (Helsinki, Finland) in collaboration with Bodil Branner (Lyngby, Denmark), Kari Hag (Trondheim, Norway), and Caroline Series (Warwick, UK)

Other Events

\underline{17:30 Urania}

Klaus Becker, Klieve, Germany From Stone–Sculpturing to Mathematics: Approximating the Inscribed Ball of Regular Polyhedra

\underline{18:00 Urania}

Movie N is a Number — a Portrait of Paul Erdős

\underline{19:30 Urania}

Phil Davis, Brown University, Providence, USA The Prospects of Mathematics in a Multimedia Civilization

\underline{20:00 Urania}

VideoMath Festival The “VideoMath Reel”, the 24 best contributions to the VideoMath competition